# Oriented Object Detection in Optical Remote Sensing Images using Deep Learning: A Survey

### v4

@@ 4.2 OBB Representations
To handle the angular boundary discontinuity, the most straightforward approach is to design a novel OBB representation. One scheme exploits new parameters to redefine the OBB, e.g., quadrilateral representation (Xu et al., 2021a), DHRec (Nie and Huang, 2023). The quadrilateral representation adopts coordinates of four vertexes to represent OBB but suffers from the inconsistency of vertex sorting 222For more detail please refer to Section B of the Appendix. between predicted OBB and ground truth (Figure 7(a)). RSDet (Qian et al., 2021, 2022) introduces a modulated loss to avoid the large regression loss caused by vertex sorting, in which the order of the four vertices of the predicted OBB is moved clockwise and counterclockwise by one place respectively. Then, taking the minimum of their regression loss. Xu et al. (2021a) proposed an effective way by gliding the vertex of the horizontal anchor on each corresponding side (Figure 7(b)). Specifically, it regresses four length ratios representing the gliding offset on each corresponding side, which can eliminate the confusion caused by vertex sorting. DHRec (Nie and Huang, 2023) encodes the OBB using double HBBs derived from the sorted horizontal and vertical coordinates (Figure 7(c)). In addition, several anchor-free methods utilize keypoints to denote an OBB, which can provide rich semantic features for oriented objects. Following the RepPoints (Yang et al., 2019b), CFA (Guo et al., 2021) and Oriented RepPoints (Li et al., 2022b) utilize the deformable convolution (Dai et al., 2017) to generate a group of reppoints (Figure 7(d)). Then, a minimum bounding rectangle for each set of predicted reppoints is calculated to obtain detection results. OSKDet (Lu et al., 2022) encodes 8 ordered points (4 vertices and 4 midpoints) to represent an OBB based on the consideration that the object has more obvious features at the vertex and edge areas (Figure 7(e)). Furthermore, an orientation-sensitive heatmap is designed to better fit the shape, allowing the model to learn the orientation and shape of the rotated object implicitly. Although the angular discontinuity has been eliminated, the redesigned OBB representations still have their limitations. Quadrilateral representation is irregular and DHRec adds extra position and oblique factors to ensure the uniqueness of the representations. Meanwhile, the keypoints rely on a complicated post-processing operator to generate a rectangular box.
@@
Therefore, there is another scheme for OBB representation that only uses additional parameters to determine the orientation of objects. A classical way is the midpoint offset representation proposed by Xie et al. (2021), in which the orientation can be inferred via the midpoint offset (Figure 7(f)). However, the midpoint offset representation usually generates a parallelogram and requires a post-processing procedure for regularization. Based on midpoint offset representation, Qiao et al. (2023) analyze the geometric relationship between the OBB and its external HBB to derive the height directly from the width and two offsets, and only use five parameters to generate the high-quality OBB directly from the horizontal anchors. QPDet (Yao et al., 2023) adopt two symmetrical offsets w.r.t the quadrant points to account for rotation and aspect ratio, and a single parameter (radius ) controls the scale (Figure 7(g)). CHPDet (Zhang et al., 2022b) define a head point to indicate the orientation but require proper annotations that specify the direction of the object head in the range of (Figure 7(h)). These strategies of orientation representation can discard angle regression, thus naturally eliminating the angular boundary discontinuity.
@@ 4.3 Discussion
As stated above, an enormous amount of research effort is committed to resolving the challenges encountered by OBB regression. Redesigning novel regression loss for mainstream -based representation empowers the detectors to solve the inconsistency problem and eliminate the confusion caused by PoA, thereby enhancing the stability of network back-propagation. Especially, Gaussian distribution based methods draw upon the trigonometric encoder and joint-optimization to achieve strong performance. On the other hand, novel OBB representation schemes can avoid angle regression, in which the completely redefined OBB representations commonly rely on complex post-processing or extra constraints, while the orientation representations provide a simple yet efficient way to determine the orientation. Nevertheless, only a handful of novel OBB representation schemes take into account the inconsistency problem.
@@ 5 Feature Representation
Robust and discriminative feature representations can significantly improve both localization and classification. As a result, the most recent improvements in detection accuracy have been via research into enhancing feature representation through innovative network architectures. In this section, we review the effort devoted to improving feature representations of oriented objects, i.e.,, rotation-invariant feature representation, and advanced feature representation.
@@ 5.1 Rotation-Invariant Feature Representations
Rotation-invariance is an essential problem when learning visual feature representations for oriented objects (Lowe, 2004; Han et al., 2021a; Yu et al., 2023). The commonly used approaches, including RRoI operators (Ma et al., 2018; Yang et al., 2018; Ding et al., 2019) and random rotation data augmentation (Han et al., 2021a), are sub-optimal, as they can only extract approximately rotation-invariant features (Lenc and Vedaldi, 2015; Worrall et al., 2017). Recently, the exploration of rotation-sensitive feature extraction networks has brought new insight to the community, which utilize different channels to represent feature information from different orientations, e.g., oriented response convolution (ORConv) (Zhou et al., 2017), group equivariant convolutional neural networks(G-CNN) (Cohen and Welling, 2016; Worrall et al., 2017; Marcos et al., 2017; Weiler and Cesa, 2019; Weiler et al., 2018). Naturally, several works, e.g., RRD (Liao et al., 2018), CHPDet (Zhang et al., 2022b), SANet (Han et al., 2022), replace ordinary convolution modules with ORConv to obtain orientation-dependent responses which are then transformed to rotation-invariant features using ORAlign and ORPooling. Besides, ReDet (Han et al., 2021a) incorporates G-CNN into the detector for rotation-equivariant feature generation, whereafter, a rotation-invariant roI align operator is designed to adaptively extract rotation-invariant features from equivariant ones according to the predicted orientations. In contrast to ORConv and G-CNN extracting features via static kernels oriented in a group of fixed orientations, ARC (Pu et al., 2023) rotates the convolution kernels dynamically according to the orientation of the objects, where the orientation is predicted in a data-dependent manner. As a result, ARC can capture the rotation information of objects with different orientations and boost the feature representation under the oriented object detection scenario.
@@ 5.2 Advanced Feature Representations
In addition to rotation-invariant networks, advanced feature extraction networks play a crucial role in high-precision detection, e.g., Vision Transformer (Vaswani et al., 2017; Dosovitskiy et al., 2021; Liu et al., 2021b), ConvNext (Liu et al., 2022d). Recently, Vision Transformer has achieved significant success in computer vision (Dosovitskiy et al., 2021; Liu et al., 2021b; Han et al., 2023), which mainly adopts the self-attention mechanism to capture global feature representations. Benefiting from the exceptional feature representation capacity, more and more methods introduce Vision Transformer for object detection and obtain impressive performance. ViT series (Dosovitskiy et al., 2021; Xu et al., 2021b; Zhang et al., 2023b) and Swin Transformer (Liu et al., 2021b) are representative architectures that can directly serve as the backbone networks for object detection, showing more excellent feature representation capability than CNN. An unsupervised pertaining scheme, MAE (He et al., 2022), has made notable progress in developing ViT for object detection. As a result, the powerful Vision Transformer architectures contribute to establishing a solid foundation for delivering outstanding achievements in oriented object detection. Nevertheless, their utilization in oriented object detection is fairly unexplored, e.g., an essential problem is how to extract rotation-related features. Based on ViT, Wang et al. (2022) design rotated varied-size window attention (RVSA) to adaptively obtain locally oriented windows at different sizes, locations, and angles, enabling detectors to extract more useful oriented information for oriented object detection tasks. Although RVSA outperforms all previous methods, it relies on the self-attention mechanism to create oriented windows, without explicitly leveraging guiding information. STD (Yu et al., 2023) adopts a controlled scheme for manipulating the feature extraction process according to the decoupled OBB parameters, i.e., center position, sizes, and angles. It follows a divide-and-conquer fashion that estimates the position, size, and angle via separate network branches at different stages. Further, the cascaded activation masks created by the decoupled OBB parameters are integrated to gradually enhance features extracted by stacked Transformer blocks. The progressive refinement of feature representation enables STD to reach state-of-the-art performance, achieving and mAP on DOTA-V1.0 (Xia et al., 2018) and HRSC2016 (Liu et al., 2016b) datasets, respectively. Recently, ConvNeXt (Liu et al., 2022d) gradually modified the standard ResNet according to a series of design decisions of Swin Transformer (Liu et al., 2021b) and demonstrated pure CNNs outperform the vision transformers in terms of accuracy and robustness. Meanwhile, ConvNeXt can maintain the efficiency of standard CNNs, thus becoming the dominant architecture in many applications.
@@ 6 Common Issues in RS scenarios
In addition to the specific challenges associated with oriented object detection, several common issues still exist regarding RS scenarios, e.g., scale variations, large aspect ratio, and complex background, as shown in Figure 8.
@@ 6.1 Scale Variations
As the ground sampling distance (GSD) can range from a few centimeters to hundreds of meters, the RS images taken by different sensors usually have large-scale variations. Additionally, while different kinds of objects may have large-scale variations, the instances of the same category also vary in size. Therefore, the inter-class and intra-class scale variations pose additional challenges. Especially, the most critical challenge mainly focuses on small objects due to their insufficient feature information, inaccurate localization, and inadequate positive samples (Cheng et al., 2023). Worse still, things get tougher when it comes to oriented objects due to the extra orientation regression and limited overlaps with anchor boxes.
@@
In recent years, a large number of effective strategies have been made to enhance the robustness and adaptability of detectors for objects with various scales, which can be classified into two categories network-level methods and data-level methods. The former committed to developing novel network structures for multi-scale features extraction, e.g., feature pyramid architectures (Lin et al., 2017) and its variants (Tian et al., 2024), and multi-branch architectures (Li et al., 2019). The latter strives to design data augmentation strategies that are independent of the network architectures and can be generalized to any detectors. Multi-scale training and testing is a useful data augmentation approach that scaling input images at different resolutions (Singh and Davis, 2018; Singh et al., 2018), has been shown to reduce overfitting and improve generalization (Ren et al., 2017; Russakovsky et al., 2015). However, it will inevitably lead to poor time efficiency.
@@
Apart from enhancing the semantic representation or amplifying objects’ size, several optimal assignment strategies have emerged to enable adequate sample assignments for small-oriented objects. Since the anchor-free methods can generate more positive samples of small objects that are apt to be ignored in the anchor-based methods, a straightforward scheme is generating samples in an anchor-free manner rather than an anchor-based manner and then refining them to obtain high-quality detection results, e.g., APE (Zhu et al., 2020), AOPG (Cheng et al., 2022a). Furthermore, Liang et al. (2022) propose a dynamic enhancement anchor network that combines the advantages of anchor-free and anchor-based methods and uses an interactive sample screening procedure to generate higher-quality training samples. In contrast to using static priors (e.g., anchor boxes or points) with fixed stride, Xu et al. (2023) design a dynamic prior with a coarse-to-fine assigner. Concretely, it first uses deformable convolution (Dai et al., 2017) to adaptively adjust the prior location and then leverages the coarse prior matching and finer posterior constraint to dynamically assign samples. Such a strategy can adaptively assign positive or negative samples according to the objects’ shape and posterior information, boosting the performance of mainstream detectors on small objects.

### v5

@@ 4.2 OBB Representations
Methods | Advantages | Disadvantages |
Quadrilateral representation [172, 125, 126] | A direct method that can compactly enclose oriented objects with large deformation and has been widely adopted to annotate objects in large-scale RS datasets. | It suffers from the inconsistency of vertex sorting, and can only represent irregular quadrilaterals, but not rectangles. |
Gliding Vertex [176] | A concise but effective method that can eliminate vertex sorting. | It can only represent irregular quadrilaterals, but not rectangles. It cannot accurately represent nearly horizontal objects. |
DHRec [116] | It can directly use horizontal object detectors to regress OBBs, thus freeing it from the trouble of PoA and vertex sorting. | It requires too many parameters to ensure the uniqueness of OBB representation. |
CFA and Oriented RepPoints [38, 40, 74] | The adaptive points learning method can capture the geometric information of oriented objects and avoid vertex sorting and PoA. | It requires complicated post-processing operations to convert reppoints into an OBB. |
OSKDet [86] | The keypoints can capture the critical features of vertex and edge area, which could better match the object shape. The unordered keypoints representation scheme can avoid the confusion of vertex sorting. | It requires complicated post-processing operations to convert irregular keypoints into an OBB. |
Midpoint Offset [174] | A concise yet effective OBB representation method that can eliminate PoA. | It generally generates a parallelogram and requires a post-processing procedure for regularization. |
QPDet [183] | A simple OBB representation method, needing just five parameters, which can avoid the generation of irregular bounding boxes. | It will suffer from new PoA of and when representing nearly horizontal objects. |
CHPDet [221] | Employ a head point to indicate the direction, effectively eliminating PoA. | The accuracy of the orientation depends on the precision of the center point and the head point. When the positions of either are inaccurate, it will affect the overall precision of the target’s location. |
To handle the angular boundary discontinuity, the most straightforward approach is to design a novel OBB representation. One common scheme exploits new parameters to redefine the OBB, e.g., quadrilateral representation [176], DHRec [116]. The quadrilateral representation adopts coordinates of four vertexes to represent OBB but suffers from the inconsistency of vertex sorting222For more detail please refer to Sec. B of the Appendix. between predictions and GT, as shown in Fig. 12a. RSDet [125, 126] mitigates this by introducing a modulated loss that considers different vertex orderings, minimizing the regression loss across these variations. Furthermore, [176] propose an effective way by gliding the vertex of the horizontal anchor on each corresponding side, as shown in Fig. 12b. Specifically, it regresses four length ratios representing the gliding offset on each corresponding side, eliminating the confusion caused by vertex sorting. To remedy the confusion for nearly horizontal objects, it directly selects horizontal detection guided by the predicted obliquity factor, albeit with slight imprecise regression. DHRec [116] encodes the OBB using double HBBs derived from the sorted horizontal and vertical coordinates, as shown in Fig. 12c. Hence, such a method allows harnessing any horizontal object detector to predict OBB.
In addition, several anchor-free methods utilize keypoints to denote an OBB, which can provide rich semantic features for oriented objects. Following the RepPoints [188], CFA [38, 40] and Oriented RepPoints [74] utilize the deformable convolution [26] to generate a group of reppoints, as shown in Fig. 12d. A minimum bounding rectangle is then computed for each set of predicted reppoints to yield detection results. OSKDet [86] encodes 8 ordered points (4 vertices and 4 midpoints) to represent an OBB based on the consideration that the object has more obvious features at the vertex and edge areas, as shown in Fig. 12e. Furthermore, an orientation-sensitive heatmap is designed to better fit the shape, allowing the model to learn the orientation and shape implicitly.
Despite the angular discontinuity has been eliminated, the redesigned OBB representations still have their limitations. Quadrilateral representation is irregular and DHRec adds extra position and oblique factors to ensure the uniqueness of the representations. Meanwhile, the keypoints rely on a complicated post-processing operator to generate a rectangular box.
@@
Therefore, there is another scheme for OBB representation that only uses extra parameters to determine the orientation of objects. The classical Midpoint Offset representation [174] infers orientation via midpoint offset, as shown in Fig. 12f, but typically produces parallelograms requiring regularization. Based on Midpoint Offset representation, [123] analyze the geometric relationship between the OBB and its external HBB to derive the height directly from the width and two offsets, only using five parameters to generate the high-quality OBB directly from the horizontal anchors. QPDet [183] adopt two symmetrical offsets w.r.t the quadrant points to account for rotation and aspect ratio, and a single parameter (radius ) controls the scale, as shown in Fig. 12g. CHPDet [221] define a head point to indicate the orientation but require proper annotations that specify the direction of the object head in the range of , as shown in Fig. 12h. These strategies of orientation representation can discard angle regression, thus naturally eliminating the angular boundary discontinuity. Additionally, the pros and cons of above OBB representation methods are summarized in Tab. 2.
@@ 4.3 Discussion
As stated above, an enormous amount of research effort is committed to resolving the challenges encountered by OBB regression. Redesigning novel regression loss for mainstream -based representation empowers the detectors to solve the inconsistency problem and eliminate the confusion caused by PoA, thereby enhancing the stability of network back-propagation. Especially, Gaussian distribution based methods draw upon the trigonometric encoder and joint optimization to achieve strong performance. On the other hand, novel OBB representation schemes can avoid orientation regression, in which the completely redefined OBB representations commonly rely on complex post-processing or extra constraints, while the orientation representations provide a simple yet efficient way to determine the orientation. Nevertheless, only a handful of novel OBB representation schemes take into account the inconsistency problem.
@@ 5 Feature Representation
Robust and discriminative feature representations play a pivotal role in improving both localization and classification. As a result, the most recent improvements in detection accuracy have been via research into enhancing feature representation through innovative network architectures. In this section, we review the effort devoted to improving feature representations of oriented objects, i.e., rotation-invariant feature representation, and advanced feature representation. Several milestone methods are shown in Fig. 13.
@@ 5.1 Rotation-Invariant Feature Representations
Rotation-invariance is an essential problem when learning visual feature representations for oriented objects [95, 48, 191]. The commonly used approaches, including RRoI operators [113, 190, 29] and random rotation data augmentation [48], are sub-optimal, as they can only extract approximately rotation-invariant features [98, 156].
Recently, the exploration of rotation-sensitive feature extraction networks brinf new insight to the community, which utilize different channels to represent feature information from different orientations, e.g., oriented response convolution (ORConv) [227], group equivariant convolutional neural networks(G-CNN) [18, 156, 114, 152, 157]. Building on this, several works, e.g., RRD [106], CHPDet [221], S2ANet [47], replace ordinary convolution modules with ORConv to obtain orientation-dependent responses which are then transformed to rotation-invariant features using ORAlign and ORPooling. Additionally, ReDet [48] incorporates G-CNN into the detector for rotation-equivariant feature generation, whereafter, a rotation-invariant roI align operator is designed to adaptively extract rotation-invariant features from equivariant ones according to the predicted orientations. Furthermore, [97] introduce selective rotation of the kernel (SKR) to enhance classification features. SKR module rotates the convolution kernel at different angles to extract rotation-invariant features, in which the output channel dimension of corresponding rotated convolution kernel is obtained by network learning adaptively.
In contrast to ORConv and G-CNN extracting features via static kernels oriented in a group of fixed orientations, ARC [121] rotates the convolution kernels dynamically according to the orientation of the objects, where the orientation is predicted in a data-dependent manner. As a result, ARC can capture the rotation information of objects with different orientations and boost the feature representation under the oriented object detection scenario.
However, [155] point out that cyclic shift phenomena exhibited by features after rotation-equivariant network are unstable due to the dilation and pooling operations. To maintain rotation-equivariant for features, a rotation-robust prototype generation (RPG) is designed, including a stabilization module and an enhancement module. The former aggregates features of different orientations to generate rotation-robust prototype, while the latter employ the prototype to enhance original features in each group.
@@ 5.2 Advanced Feature Representations
In addition to rotation-invariant networks, advanced feature extraction networks, particularly Vision Transformer [151, 22, 85], have played a crucial role in high-precision detection. ConvNeXt [94] is another notable architecture that has also contributed to the field, though its discussion will be brief in this context.
Recently, Vision Transformer has achieved significant success in computer vision [22, 85, 63], primarily due to its self-attention mechanism that captures global feature representations. This exceptional capacity for feature representation lead to its increasing adoption in object detection, yielding remarkable results. Representative architectures like the ViT series [22, 181, 224] and Swin Transformer [85] can serve directly as the backbone networks, showing more excellent feature representation capability than CNN. Furthermore, the unsupervised pertaining scheme, MAE [46], makes notable progress in developing ViT for object detection. Consequently, the powerful Vision Transformer architectures contribute to establishing a solid foundation for delivering outstanding achievements in oriented object detection.
Nevertheless, their utilization in oriented object detection is fairly unexplored, e.g., an essential problem is how to extract rotation-related features. [169] addressed this by designing rotated varied-size window attention (RVSA) based on ViT, which adaptively generates locally oriented windows at different sizes, locations, and angles. Although RVSA outperforms all previous methods, it relies on the self-attention mechanism to create oriented windows, without explicitly leveraging guiding information.
STD [191], on the other hand, adopts a controlled scheme for manipulating the feature extraction process according to the decoupled OBB parameters, i.e., center position, sizes, and angles. It follows a divide-and-conquer fashion that estimates the position, size, and angle via separate network branches at different stages. Further, the cascaded activation masks created by the decoupled OBB parameters are integrated to gradually enhance features extracted by stacked Transformer blocks. The progressive refinement of feature representation enables STD to reach state-of-the-art performance, achieving and mAP on DOTA-V1.0 [172] and HRSC2016 [100] datasets, respectively.
More recently, ConvNeXt [94] gradually modified the standard ResNet according to a series of design decisions of Swin Transformer [85] and demonstrated pure CNNs outperform the vision transformers in terms of accuracy and robustness. Meanwhile, ConvNeXt can maintain the efficiency of standard CNNs, thus becoming the dominant architecture in many applications.
@@ 6 Common Issues in RS scenarios
In addition to the specific challenges associated with oriented object detection, several common issues still exist regarding RS scenarios, e.g., complex background, scale variations, large aspect ratio, and lack of annotated samples, as shown in Fig. 14. Several representative methods for tracking these common issues are shown in Fig. 15.
6.1 Complex Backgrounds
Due to the wide visual field and complex earth’s surface, RS images typically contain a variety of complex backgrounds, causing significant interference in detection. Objects are frequently surrounded by different backgrounds, necessitating detectors with heightened discriminative capabilities. Besides, the presence of backgrounds with textures and shapes resembling the objects leads to a high incidence of false positives.
To this end, a series of efforts have been made to suppress background noise and emphasize the valuable areas of the objects. CFC-Net [112] and PETDet [107] both combine channel attention and spatial attention modules to learn the semantic correlation between foreground and background. However, these methods rely on a self-attention mechanism and lack direct foreground guidance. To enhance the discrimination of foreground, [236] proposes a foreground relation module for foreground-contextual representations under the supervision of the designed foreground map. Similarly, CBDA-Net [105] builds two parallel spatial attention streams to capture center and boundary attention features, which could assist detectors in improving object localization accuracy. Besides, SCRDet [199] adopts the pixel attention network to generate a saliency map that can separate foreground from background, and use Squeeze-and-Excitation (SE) blocks [61] as the channel attention network to enhance the saliency map further.
Building on SCRDet, SCRDet++ [194] proposes an instance-level denoising (InLD) module that weakens the feature response of the background region while decouples the features of different categories into their corresponding channels. Similar to InLD, several methods [229, 197, 182, 219] introduce semantic mask modules that is supervised by the class-wise mask transformed from the oriented GT. The semantic mask modules separate the features of different categories in channel dimension, which can facilitate the networks to reduce both the background and interclass interference.
@@ 6.2 Scale Variations
As the ground sampling distance (GSD) can range from a few centimeters to hundreds of meters, the RS images taken by different sensors usually have large-scale variations. Additionally, even within the same category, object instances exhibit wide-ranging sizes. These inter-class and intra-class scale variations pose additional challenges. Especially, the most critical challenge mainly focuses on small objects due to their insufficient feature information, inaccurate localization, and inadequate positive samples [21]. Worse still, things get tougher when it comes to oriented objects due to the extra orientation regression and limited overlaps with anchor boxes.
@@
In recent years, a large number of effective strategies have been made to enhance the robustness and adaptability of detectors for objects with various scales, which can be classified into two categories: network-level and data-level methods. Network-level methods are committed to developing novel network structures for multi-scale features extraction, e.g., feature pyramid architectures [77] and its variants [148], and multi-branch architectures [75, 120, 11]. [120] design a feature selection module to adjust receptive fields, which proposes a channel attention network to adaptively fuse the feature extracted by using kernels of various sizes, aspect ratios, and orientations. PKINet [11] employs parallel multi-scale convolution kernels without dilation to effectively capture features across varying receptive fields. Data-level methods, on the other hand, strive to design data augmentation strategies that are independent of the network architectures and can be generalized to any detectors. Multi-scale training and testing is a useful data augmentation approach that scaling input images at different resolutions [137, 141], has been shown to reduce overfitting and improve generalization [130, 131, 128]. However, it will inevitably lead to poor time efficiency.
@@
Beyond enhancing the semantic representation or amplifying object size, several optimal assignment strategies have emerged to enable adequate sample assignments for small-oriented objects. Since the anchor-free methods can generate more positive samples of small objects that are apt to be ignored in the anchor-based methods. Consequently, approaches like APE [208] and AOPG [19] generate samples in an anchor-free manner, refining them for high-quality detection results. [80] propose a dynamic enhancement anchor network that combines the advantages of anchor-free and anchor-based methods and uses an interactive sample screening procedure to yield higher-quality training samples.
Rather than relying on IoU alone, several methods [115, 223, 31, 55] collaborate the prior (e.g., IoU between anchor and GT) and posterior information (e.g., classification and localization confidence) as the evaluation criteria to screen positive samples. Such a scheme can push the detectors to dynamically mine high quality anchors as positive. In contrast to using static priors (e.g., anchor boxes or points) with fixed stride, [175] design a dynamic prior with a coarse-to-fine assigner. Concretely, it first uses deformable convolution [26] to adaptively adjust the prior location and then leverages the coarse prior matching and finer posterior constraint to dynamically assign samples. Such a strategy can adaptively assign positive or negative samples according to the objects’ shape and posterior information, boosting the performance of mainstream detectors on small objects.